I have this problem. Kirsten is 4 times Claires age. Emily is 6yrs older than Claire and 6yrs younger than Zack. if anna is twice as old as Claire and the total of their ages is 81, How hod is Kirsten? I can't even get my equation started.
Answer by richwmiller(17219)

I'll get you going.
k=4c
e=c+6
e=z-6
a=2c
k+c+e+z+a=81
5 equations 5 unknowns
Have fun. Be sure to set c+6=z-6 and solve for z
then using only c's solve for c in the long equation